Trigger Operation Meeting (9/4/07)
Attendees: Debbie Bard, Kim Hojeong, Su Dong, Selina Li, Mark Tibbetts.
Date : Tuesday, September 4th, 2007
Time : 2:00pm - 3:30pm PST
SLAC Room : Madrone conf room, ROB
Dial-in : 1-510-665-5437, pass code 6649#
--- Minutes ---
- During weekends, there were order of ten 100% deadtime due to EMC/EMT/DIRC/ODF. All cleared by starting a new run.
- We planned to power off all the crate today after operations meeting, but we changed __not__ to power off for a while, since IFR people want to take cosmics for next few weeks.
- Run 76397 has warm EMT tower not associated with hot EMC tower (e-logbook entry, look for [03:35]). It disappeared in new run and never came back. The cause is unclear, maybe EMC UPC, maybe TPB? We will keep an eye on it when we start to take data.
- Before Run7 begins, we need to check if newest tag is being used.
- Arik's Track fixup validation is still on-going. tcstage is not working currently.
- Su Dong and Selina will do some R&D work on L3 EMT filter.
Trickle injection monitoring
- At the end of Run7, they plan to burn the detectors down to acheive highest delivered luminosity. We need to prepare for factor 2~3 trigger rate. Current emergency configuration which reduces 35% is not enough.
- We will have another meeting next week, and then we'll go into every other week mode.
- Rainer prepared a plot, but we forgot to call in! Sorry Rainer.
- Many thanks to Debbie for her hard work for a year!
This page was created by Kim Hojeong (hojeong@slac); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.
Last Modified: September 4, 2007.